OpenStudy (anonymous):

While driving on the freeway at 110 km h, you pass a truck whose total length you estimate at 25 m. If it takes you, in the driver’s seat, 5.5 s to pass from the rear of the truck to its front, what is the truck’s speed relative to the road?

OpenStudy (anonymous):

Normally, it would take 9/11 seconds to pass the truck if it is stationary. From that, I got it takes the driver 4.68 additional seconds to pass it than if it was stationary. 110 = other_rate + 4.68 (This isn't right, and I don't think it is, but it is the closest equation I can think of.)

OpenStudy (ujjwal):

If the answer is 26 m/s or 93.6 Km/h .. I can help you!

OpenStudy (ujjwal):

@DanielHendrycks

OpenStudy (ujjwal):

Do you have the answer?

OpenStudy (ujjwal):

Let the velocity of truck relative to ground be v. Find the relative velocity. Displacement= relative velocity \(\times\) time.. Convert Km/h into m/s since you have all other values in SI.

OpenStudy (anonymous):

Sorry. I'm here. Give me a second to read.

OpenStudy (anonymous):

It is 26 m/s

OpenStudy (ujjwal):

Ok then do as i have mentioned!

OpenStudy (anonymous):

The truck isn't going to make and distance from the ground (it will keep touching it). Do you mean the point we should be speaking relative to is a place on the ground?

OpenStudy (ujjwal):

I don't get what you are trying to say!

OpenStudy (anonymous):

You are saying it should be relative to the ground right?

OpenStudy (ujjwal):

Ok i am explaining it all again:

OpenStudy (ujjwal):

Let the velocity of truck relative to ground be v. Now, find the velocity of truck relative to you.. Now, Displacement of truck relative to you= velocity of truck relative to you \(\times\) time In this case, the displacement of truck relative to you is 25 m.

OpenStudy (anonymous):

25 = v relative to me * 5.5; therefore v relative to me is 25/5.5 or 4.54. That's not right.

OpenStudy (ujjwal):

Velocity of truck relative to you=your velocity-velocity of truck relative to ground= (30.55-v) m/s Note: 110 Km/h=30.55 m/s Now, 25=(30.55-v)* 5.5 Or, v=26 m/s

OpenStudy (anonymous):

Got it! I'll make note of this problem.

OpenStudy (anonymous):

RAPE JOKE